OpenStudy (dmezzullo):

Select the equations that are parallel and perpendicular to y = 2x - 1 and that pass through the point (4, 1). parallel: y = 2x - 7 perpendicular: y = negative one over twox + 3 parallel: y = 2x + 2 perpendicular: y = negative one over twox + four and one half parallel: y = negative one over twox - 7 perpendicular: y = 2x + 3 parallel: y = -2x - 7 perpendicular: y = negative one over twox - 1

OpenStudy (anonymous):

1) parallel lines have equal slopes 2) perpendicular lines have slopes that are negative reciprocals of eachother.

OpenStudy (dmezzullo):

I don't understand wut u mean

OpenStudy (anonymous):

let parallel line be \(y=mx+b\) then \(m=2\) find "b" using the fact that it passes through (4,1)

OpenStudy (anonymous):

**PROPERTY: if two lines are parallel: then their slopes are equal.

OpenStudy (anonymous):

In case of parallel lines slopes are equal..

OpenStudy (dmezzullo):

ok

OpenStudy (anonymous):

so, what did you gt for "b"?

OpenStudy (anonymous):

Now, parallel lines that is given has slope of 2.. So use this formula: \[y - y_1 = m(x - x_1)\]

OpenStudy (dmezzullo):

(2,1)

OpenStudy (anonymous):

"b" is just a number -> the y-intercept.

OpenStudy (anonymous):

Here : \[y - 1 = 2(x - 4)\] Solve this..

OpenStudy (dmezzullo):

y-1=2x-8

OpenStudy (dmezzullo):

+1 on both sides y=2x+9

OpenStudy (anonymous):

On LHS, you should have y only..

OpenStudy (anonymous):

Yep.. Follow the same procedure now for second one.. But remember: In case of perpendicular lines: slopes are negative reciprocal of each other..

OpenStudy (anonymous):

2x - 7 should be there..

OpenStudy (anonymous):

y - 1 = 2x - 8 By adding 1 to both the sides what will you get??

OpenStudy (anonymous):

Now find the slope of given equation in case of perpendicularity..

OpenStudy (dmezzullo):

idk wut u mean

OpenStudy (anonymous):

How it is 2x - 9 ??

OpenStudy (anonymous):

y - 1 = 2x - 8 Now add 1.. y - 1 + 1 = 2x - 8 + 1 ??? How much??

OpenStudy (dmezzullo):

y=2-9

OpenStudy (anonymous):

Wait, now you are doing more wrong.. -8 + 1 = ???

OpenStudy (dmezzullo):

Oh -7

OpenStudy (anonymous):

Yep..

OpenStudy (anonymous):

Simply find the slope of other line you are given with..

OpenStudy (anonymous):

Oh My mistake, no other line is given.. Now you know that the slope is 2 for the line you are given.. So to find the line perpendicular to it, take negative reciprocal of the slope.. So, Negative Reciprocal of 2 is ??

OpenStudy (anonymous):

See if you have two lines, one has slope m1 and second has slope m2, then: \[m_1 m_2 = -1\] If m1 = 2 here then what will be m2?? can you tell??

OpenStudy (anonymous):

This is true only in case of perpendicular lines, I forgot to mention it above..

OpenStudy (dmezzullo):

I'm confued

OpenStudy (dmezzullo):

confused*

OpenStudy (anonymous):

parrallel lines have same slope different intercepts while perpendicular lines have opposite reciprocal slope

OpenStudy (anonymous):

Let me explain it.. 

OpenStudy (anonymous):

Suppose these are the two lines that are perpendicular to each other.. Suppose one line has slope \(m_1\) and other line has slope of \(m_2\)., then there is a relation between m1 and m2, that is : \(m_1m_2 = -1\).. Getting this much?

OpenStudy (anonymous):

@dmezzullo getting or not??

OpenStudy (dmezzullo):

Not so much

OpenStudy (dmezzullo):

sorry had to take my puppy out

OpenStudy (anonymous):

Where are you having problem?

OpenStudy (dmezzullo):

Jus understanding it all